OpenStudy (anonymous):

4z1over 6=5z+7 over 6

OpenStudy (saifoo.khan):

\[\frac{4z}{6}=\frac{5z+7}{6} \]

OpenStudy (saifoo.khan):

\[\frac{4z}{\cancel6}=\frac{5z+7}{\cancel{6} }\]

OpenStudy (saifoo.khan):

finally, 4z = 5z + 7 Solve. :)

OpenStudy (diyadiya):

-7